# Carleman Estimates and Controllability of Forward Stochastic Parabolic Equations with General Dynamic Boundary Conditions

### v1

@@ 1 Introduction and Main Results
The equation (1.7) with $u\equiv v_{1}\equiv v_{2}\equiv 0$ models diffusion phenomena, such as the evolution of quantities like heat, bacterial populations, or chemical concentrations, under stochastic perturbations and interactions with the boundary through the dynamic boundary conditions ((1.7)2 and (1.7)3), which involve the time derivative of the state on the boundary. Additionally, the convection terms $B_{1}\cdot\nabla y$ and $B_{2}\cdot\nabla_{\Gamma}y_{\Gamma}$ in (1.7) enable the modeling of these diffusion processes in a moving medium, such as fluid flow with a specific velocity. For further details on the model of stochastic diffusion equations, we refer to [10, 27] and the references therein.
@@
The second main result of this paper is the following approximate controllability of (1.7).
@@
Equation (1.7) is approximately controllable at time $T$ and for any control region $G_{0}\Subset G$. That is, for any initial state $(y_{0},y_{\Gamma,0})\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$ and final state $(y_{T},y_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$, and for all $\varepsilon>0$, there exist controls
$(\widehat{u},\widehat{v}_{1},\widehat{v}_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma)),$
such that the corresponding solution $(\widehat{y},\widehat{y}_{\Gamma})$ to (1.7) satisfies
$\mathbb{E}\|(\widehat{y}(T,\cdot),\widehat{y}_{\Gamma}(T,\cdot))-(y_{T},y_{\Gamma,T})\|^{2}_{\mathbb{L}^{2}}\leq\varepsilon.$
The third main result is the following null controllability of (1.7).
@@
Equation (1.7) is null controllable at time $T$ and for any control region $G_{0}\Subset G$. That is, for any initial state $(y_{0},y_{\Gamma,0})\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$, there exist controls
$(\widehat{u},\widehat{v}_{1},\widehat{v}_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma)),$
such that the associated solution $(\widehat{y},\widehat{y}_{\Gamma})$ to (1.7) satisfies
$(\widehat{y}(T,\cdot),\widehat{y}_{\Gamma}(T,\cdot))=(0,0)\quad\textnormal{in}\;\,G\times\Gamma,\quad\textnormal{a.s.}$
Moreover, the controls $\widehat{u}$, $\widehat{v}_{1}$, and $\widehat{v}_{2}$ can be chosen such that
$\begin{aligned} \|\widehat{u}\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))}+\|\widehat{v}_{1}\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G))}+\|\widehat{v}_{2}\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))}\leq e^{CK}\|(y_{0},y_{\Gamma,0})\|^{2}_{L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})},\end{aligned}$ (1.8)
where $K$ has the following form:
Remark 1.2.
In (1.7), by acting on the system with the controls $u$, $v_{1}$, and $v_{2}$, we aim to influence the final behavior of the evolution. The conormal derivative $\partial_{\nu}^{\mathcal{A}}y$ serves as the coupling term between the bulk equation in $Q$ and the surface equation on $\Sigma$. Thus, the controls $u$, $v_{1}$, and $v_{2}$ directly affect the drift term through the control $u$ on the small region $G_{0}$ of the bulk equation, and the diffusion terms of both the bulk and surface equations via the extra controls $v_{1}$ and $v_{2}$, which act throughout the entire spatial domain $G$. Through this coupling term $\partial_{\nu}^{\mathcal{A}}y$, these controls also influence the drift term of the surface equation. This interplay allows us to establish the null and approximate controllability of (1.7).
@@
This paper investigates the null and approximate controllability of a class of stochastic parabolic equations with dynamic boundary conditions. For results on the controllability of deterministic parabolic equations with such boundary conditions, we refer the reader to [28], [1], and [20]. The case of fully discrete parabolic equations is addressed in [21]. Additionally, numerous results on the controllability of stochastic partial differential equations are available in the literature; for a comprehensive discussion, we refer the reader to the book [27] and the references therein.
@@
The controllability of stochastic parabolic equations with various boundary conditions—both static and dynamic—has been extensively studied. Specifically, for static boundary conditions, controllability has been investigated for Dirichlet, Neumann, and Robin boundary conditions. For Dirichlet boundary conditions, [2] was among the first to establish controllability results for forward stochastic parabolic equations with a localized control in the drift term, under certain restrictive assumptions on the system’s coefficients. Later, [32] extended these results using global Carleman estimates to address the controllability of general forward stochastic parabolic equations, introducing an additional control in the diffusion term and considering only bounded adapted zero-order coefficients that may depend on space, time, and random variables. We also refer to [14] where a general weighted identity for some stochastic operators with complex principal parts is established and then used to study controllability of certain type of stochastic partial differential equations as stochastic complex Ginzburg–Landau equations. In [32], the authors achieved these results by introducing a second control acting on the diffusion term across the entire domain. The work [3] further investigates the controllability of more general stochastic parabolic equations, incorporating both zero- and first-order terms, under the assumption that all coefficients are only bounded and adapted. In [24, 35], the Lebeau-Robiano strategy was applied to explore both approximate and null controllability of the forward stochastic heat equation with a single control in the drift term, assuming space-independent coefficients. Additionally, the controllability of some classes of stochastic parabolic equations with Robin boundary conditions has been studied in [7, 34]. For results on the null controllability of a class of stochastic semidiscrete parabolic equations, see [37].
@@
For controllability results of stochastic parabolic equations with dynamic boundary conditions, we refer to [6] for forward and backward stochastic parabolic equations with only reaction terms, [5] for a class of one-dimensional stochastic heat equations, [4] for backward stochastic parabolic equations that incorporate both reaction and convection terms, and [8] for the case of the forward stochastic heat equation with a time-space localized control in the drift term of the bulk equation, without extra forces. For controllability results concerning certain coupled forward-backward stochastic parabolic equations, we refer the reader to [23, 29].
@@
Following the classical duality argument, the controllability of linear stochastic partial differential equations, as explored in the previously cited works, can be proven through an observability problem for the corresponding adjoint equation. This involves unique continuation for approximate controllability and an observability inequality for null controllability. A key tool in proving these observability results (for both parabolic PDEs and SPDEs) is Carleman estimates, which provide energy estimates with exponential weights and parameters sufficiently large to absorb undesired lower-order terms. Carleman estimates, first introduced by Carleman in 1939 to prove the uniqueness of solutions for second-order elliptic partial differential equations with two variables ([9]), are essential not only for analyzing control problems (see [27]) but also for solving certain inverse problems (e.g., [25, 26]).
@@
To the best of our knowledge, this paper is the first to investigate the null and approximate controllability of a more general class of linear forward stochastic parabolic equations with dynamic boundary conditions. This class includes both reaction and convection terms, as well as general second-order parabolic operators. Furthermore, we establish the estimate (1.8) for the null-controllability cost of (1.7), which is expressed in terms of the final control time $T$ and the coefficients $a_{1}$, $a_{2}$, $B_{1}$, and $B_{2}$.
@@
In equation (1.7), since the controls $v_{1}$ and $v_{2}$ influence the diffusion terms in the whole domain $G$ and on its boundary $\Gamma$, respectively, the controllability results derived for (1.7) can be generalized to cases involving the state $(y,y_{\Gamma})$ and the gradients $\nabla y$ and $\nabla_{\Gamma}y_{\Gamma}$ in the diffusion terms. Specifically, the diffusion terms $v_{1}\,dW(t)$ (resp., $v_{2}\,dW(t)$) can be replaced by the more general expressions $b_{1}y+\widetilde{B}_{1}\cdot\nabla y+v_{1}]\,dW(t)$ (resp., $b_{2}y_{\Gamma}+\widetilde{B}_{2}\cdot\nabla_{\Gamma}y_{\Gamma}+v_{2}]\,dW(t)$), where the coefficients $b_{i}$ and $\widetilde{B}_{i}$, for $i=1,2$, are adapted and bounded stochastic processes. Indeed, in the case where (1.7) is null (resp., approximately) controllable, one can find the control triple $(u,v_{1},v_{2})$ such that
@@
It then follows that the new control triple $(u^{\star},v^{\star}_{1},v^{\star}_{2})$, defined by
$u^{\star}=u,\qquad v^{\star}_{1}=v_{1}-(b_{1}y+\widetilde{B}_{1}\cdot\nabla y),\qquad v^{\star}_{2}=v_{2}-(b_{2}y_{\Gamma}+\widetilde{B}_{2}\cdot\nabla_{\Gamma}y_{\Gamma}),$
achieves the null (resp., approximate) controllability for the following extended system:
$\begin{cases}\begin{array}[]{ll}dy-\nabla\cdot(\mathcal{A}\nabla y)\,dt=(a_{1}y+B_{1}\cdot\nabla y+\mathbbm{1}_{G_{0}}u^{\star})\,dt+(b_{1}y+\widetilde{B}_{1}\cdot\nabla y+v_{1}^{\star})\,dW(t)&\textnormal{in}\,\,Q,\\ dy_{\Gamma}-\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma})\,dt+\partial_{\nu}^{\mathcal{A}}y\,dt=(a_{2}y_{\Gamma}+B_{2}\cdot\nabla_{\Gamma}y_{\Gamma})\,dt+(b_{2}y_{\Gamma}+\widetilde{B}_{2}\cdot\nabla_{\Gamma}y_{\Gamma}+v_{2}^{\star})\,dW(t)&\textnormal{on}\,\,\Sigma,\\ y_{\Gamma}=y|_{\Gamma}&\textnormal{on}\,\,\Sigma,\\ (y,y_{\Gamma})|_{t=0}=(y_{0},y_{\Gamma,0})&\textnormal{in}\,\,G\times\Gamma.\end{array}\end{cases}$

### v2

@@ 1 Introduction and Main Results
The system (1.9) with $\xi\equiv\xi_{\Gamma}\equiv u\equiv v_{1}\equiv v_{2}\equiv 0$ models diffusion phenomena, such as the evolution of heat, bacterial populations, or chemical concentrations, under stochastic perturbations and in interaction with the boundary through the dynamic boundary conditions (1.9)2–(1.9)3, which involve the time derivative of the boundary state. The convection terms $B_{1}\cdot\nabla y$ and $B_{2}\cdot\nabla_{\Gamma}y_{\Gamma}$ describe the transport of these quantities in a moving medium, such as a fluid flow with a prescribed velocity field. For further details on stochastic diffusion models, we refer to [15, 34] and the references therein.
We define the following energy functional, referred to as the sentinel, associated with equation (1.9) and localized on the bulk–surface observation sets $\mathcal{O}\subset G$ and $\mathcal{O}_{\Gamma}^{1},\mathcal{O}_{\Gamma}^{2}\subset\Gamma$:
$\begin{aligned} \mathcal{E}(y,y_{\Gamma}):=&\;\frac{1}{2}\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}}|y|^{2}\,dx\,dt+\frac{1}{2}\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}^{1}_{\Gamma}}|y_{\Gamma}|^{2}\,d\sigma\,dt\\ &\;+\frac{1}{2}\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}^{2}_{\Gamma}}|\nabla_{\Gamma}y_{\Gamma}|^{2}\,d\sigma\,dt,\end{aligned}$ (1.10)
where $(y,y_{\Gamma})$ denotes the solution of (1.9) corresponding to the parameters $\tau_{1},\tau_{2}$ and to the controls $u$, $v_{1}$, and $v_{2}$.
Introduced by J.-L. Lions (see [28]), the notion of insensitizing control consists in finding controls $(u,v_{1},v_{2})$ such that the energy functional $\mathcal{E}$ is locally invariant with respect to small perturbations of the initial data. More precisely, the insensitizing control problem associated with system (1.9) is defined as follows.
Definition 1.1.
Let $\xi\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G))$, $\xi_{\Gamma}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$, and $(y_{0},y_{\Gamma,0})\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$. A control triple $(u,v_{1},v_{2})$ is said to insensitize the functional $\mathcal{E}$ if
$\frac{\partial\mathcal{E}(y,y_{\Gamma})}{\partial\tau_{1}}\bigg|_{\tau_{1}=\tau_{2}=0}=\frac{\partial\mathcal{E}(y,y_{\Gamma})}{\partial\tau_{2}}\bigg|_{\tau_{1}=\tau_{2}=0}=0,\qquad\forall\,(\widehat{y}_{0},\widehat{y}_{\Gamma,0})\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})\ \;\textnormal{with}\;\ \mathbb{E}\|(\widehat{y}_{0},\widehat{y}_{\Gamma,0})\|_{\mathbb{L}^{2}}^{2}=1.$ (1.11)
For the sake of clarity, the following notations will be used in certain parts of this paper:
  • Controlled equation: For the null controllability problem, we denote by (CE) the system (1.9) with $\tau_{1}\equiv\tau_{2}\equiv 0$ and $\xi\equiv\xi_{\Gamma}\equiv 0$, namely,
    $\begin{cases}\begin{array}[]{ll}dy-\nabla\cdot(\mathcal{A}\nabla y)\,dt=\bigl(a_{1}y+B_{1}\cdot\nabla y+\mathbbm{1}_{G_{0}}u\bigr)\,dt+v_{1}\,dW(t)&\textnormal{in }Q,\\[3.00003pt] dy_{\Gamma}-\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma})\,dt+\partial_{\nu}^{\mathcal{A}}y\,dt=\bigl(a_{2}y_{\Gamma}+B_{2}\cdot\nabla_{\Gamma}y_{\Gamma}\bigr)\,dt+v_{2}\,dW(t)&\textnormal{on }\Sigma,\\[3.00003pt] y_{\Gamma}=y|_{\Gamma}&\textnormal{on }\Sigma,\\[3.00003pt] (y,y_{\Gamma})|_{t=0}=(y_{0},y_{\Gamma,0})&\textnormal{in }G\times\Gamma.\end{array}\end{cases}$ (CE)
  • State shorthand: For convenience, we set $Y_{0}:=(y_{0},y_{\Gamma,0})$ to denote the pair consisting of the bulk and boundary components of the state at the initial time $t=0$.
  • Control space: The admissible control space associated with system (1.9) is defined as
    $\mathcal{U}_{T}:=L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma)),$
    so that any control triple $(u,v_{1},v_{2})\in\mathcal{U}_{T}$ corresponds respectively to the distributed control in the interior, the stochastic bulk control, and the stochastic boundary control.
  • Control energy: For a control triple $(u,v_{1},v_{2})\in\mathcal{U}_{T}$, we define the associated energy by
    $E(u,v_{1},v_{2}):=\frac{1}{2}\mathbb{E}\int_{0}^{T}\Big\{\|u\|_{L^{2}(G_{0})}^{2}+\|v_{1}\|_{L^{2}(G)}^{2}+\|v_{2}\|_{L^{2}(\Gamma)}^{2}\Big\}\,dt.$
@@
The second main result of this paper concerns the null controllability of the system (CE).
@@
The system (CE) is null controllable at time $T$ for any control region $G_{0}\Subset G$. More precisely, for any initial state $Y_{0}\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$, there exists a control triple $(\widehat{u},\widehat{v}_{1},\widehat{v}_{2})\in\mathcal{U}_{T}$ such that the corresponding solution $(\widehat{y},\widehat{y}_{\Gamma})$ of (CE) satisfies
$(\widehat{y}(T,\cdot),\widehat{y}_{\Gamma}(T,\cdot))=(0,0)\quad\textnormal{in }G\times\Gamma,\quad\textnormal{a.s.}$ (1.12)
Moreover, the controls can be chosen to satisfy
$\begin{aligned} E(\widehat{u},\widehat{v}_{1},\widehat{v}_{2})\leq\exp(CK_{T})\,\mathbb{E}\|Y_{0}\|^{2}_{\mathbb{L}^{2}},\end{aligned}$ (1.13)
where the constant $K_{T}$ is given by
$K_{T}:=1+\frac{1}{T}+\|a_{1}\|_{\infty}^{2/3}+\|a_{2}\|_{\infty}^{2/3}+T(\|a_{1}\|_{\infty}+\|a_{2}\|_{\infty})+(1+T)(\|B_{1}\|_{\infty}^{2}+\|B_{2}\|_{\infty}^{2}).$ (1.14)
Once Theorem 1.2 is established, the set of admissible internal controls
[…]
is nonempty. We then define the cost of null controllability, that is, the minimal $L^{2}$-norm of the controls required to steer the system to zero at time $T$, by
$\mathcal{K}(Y_{0},G_{0}):=\inf_{(u,v_{1},v_{2})\in\mathcal{A}_{ad}(Y_{0},G_{0})}E(u,v_{1},v_{2}).$
As a direct consequence of Theorem 1.2, one can deduce the following explicit bound on the null controllability cost of the system (CE) .
Corollary 1.1.
Let $Y_{0}\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$, $T>0$, and let the coefficients $a_{i}\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(G))$ and $B_{i}\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(G;\mathbb{R}^{N}))$, $i=1,2$. Then, the cost of null controllability of (CE) satisfies
$\mathcal{K}(Y_{0},G_{0})\leq\exp(CK_{T})\,\mathbb{E}\|Y_{0}\|^{2}_{\mathbb{L}^{2}},$ (1.15)
where the constants $C$ and $K_{T}$ are the same as defined in (1.13)-(1.14).
The third main result of this paper concerns the existence of insensitizing controls for (1.9).
@@
Assume that $G_{0}\cap\mathcal{O}\neq\emptyset$, and let $\mathcal{O}_{\Gamma}^{1},\mathcal{O}_{\Gamma}^{2}\subset\Gamma$ be arbitrary nonempty open subsets, and set $(y_{0},y_{\Gamma,0})=(0,0)$. Then there exist constants $K>0$ and $C>0$, depending only on $G$, $G_{0}$, $\mathcal{O}$, $\mathcal{O}^{1}_{\Gamma}$, $\mathcal{O}^{2}_{\Gamma}$, $T$, $\beta_{0}$, $M$, $M_{\Gamma}$, and the coefficients $a_{i}$, $B_{i}$ ($i=1,2$), such that for any $\xi\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G))$, and $\xi_{\Gamma}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$, satisfying
$\mathbb{E}\iint_{Q}\exp(Kt^{-1})|\xi|^{2}\,dx\,dt<\infty,\quad\mathbb{E}\iint_{\Sigma}\exp(Kt^{-1})|\xi_{\Gamma}|^{2}\,d\sigma\,dt<\infty,$ (1.16)
there exists a control triple $(u,v_{1},v_{2})\in\mathcal{U}_{T}$ that insensitizes the functional $\mathcal{E}$ in the sense of Definition 1.1. Moreover, the controls satisfy the estimate
$E(u,v_{1},v_{2})\leq C\bigg[\mathbb{E}\iint_{Q}\exp(Kt^{-1})|\xi|^{2}\,dx\,dt+\mathbb{E}\iint_{\Sigma}\exp(Kt^{-1})|\xi_{\Gamma}|^{2}\,d\sigma\,dt\bigg].$ (1.17)
[…]
$(z,Z,\widehat{Z})=(0,0,0)\,\,\,\textnormal{in}\,\,\,Q_{0}\times Q\times\Sigma,\;\,\textnormal{a.s.}\Longrightarrow\;\,(z_{T},z_{\Gamma,T})=(0,0),\,\,\,\textnormal{in}\,\,\,G\times\Gamma,\;\,\textnormal{a.s.}$
Hence, using the above uniqueness together with a duality argument, we deduce the following approximate controllability of (CE).
Corollary 1.2.
@@
Equation (CE) is approximately controllable at time $T$ for any control region $G_{0}\Subset G$. More precisely, for any initial state $Y_{0}=(y_{0},y_{\Gamma,0})\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$, any target state $Y_{T}=(y_{T},y_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$, and any $\varepsilon>0$, there exists a control triple $(\widehat{u},\widehat{v}_{1},\widehat{v}_{2})\in\mathcal{U}_{T}$ such that the corresponding solution $(\widehat{y},\widehat{y}_{\Gamma})$ to (CE) satisfies
$\mathbb{E}\|\widehat{y}(T)-y_{T}\|^{2}_{L^{2}(G)}+\mathbb{E}\|\widehat{y}_{\Gamma}(T)-y_{\Gamma,T}\|^{2}_{L^{2}(\Gamma)}\leq\varepsilon.$ (1.18)
@@
This paper investigates the null controllability and an insensitizing control problem for a class of forward anisotropic stochastic parabolic equations with general dynamic boundary conditions. For controllability results concerning deterministic parabolic equations with such boundary conditions, we refer the reader to [1], [25], and [36]. The case of fully discrete parabolic equations is treated in [26]. For insensitizing control problems of such equations, we refer to [8, 47], and to [37], where insensitizing control is studied using a sentinel that also involves the tangential gradient. In addition, the literature on controllability for deterministic partial differential equations is vast, and it is not possible to provide an exhaustive list here; for a comprehensive account of this topic, we refer the reader to the monographs [14, 46] and the references therein.
@@
The controllability of stochastic parabolic equations with various static boundary conditions has been extensively studied. In particular, controllability results have been established for Dirichlet, Neumann, and Robin boundary conditions. For the case of Dirichlet boundary conditions, [2] was among the first works to obtain partial controllability results for forward stochastic parabolic equations with a localized control acting in the drift term, under rather restrictive assumptions on the system coefficients. Subsequently, [40] substantially extended these results by means of global Carleman estimates, addressing the controllability of more general forward stochastic parabolic equations. In that work, an additional control acting on the diffusion term was introduced, and the analysis was carried out under the assumption that the zero-order coefficients are bounded, adapted, and may depend on space, time, and random variables. We also refer to [18], where a general weighted identity for certain stochastic operators with complex principal parts is derived including stochastic complex Ginzburg–Landau operators, stochastic heat operators, and stochastic Schrödinger operators, and subsequently applied to study the controllability of several classes of stochastic partial differential equations. The work [4] further investigates the controllability of more general stochastic parabolic equations involving both zero- and first-order terms, under the assumption that all coefficients are bounded and adapted. In [4, 40], controllability results are obtained by introducing a second control acting on the diffusion term over the entire spatial domain. In contrast, in [31, 45], the Lebeau–Robbiano strategy is employed to study both approximate and null controllability of the forward stochastic heat equation with a single control acting in the drift term, under the assumption of space-independent coefficients. In addition, the controllability of certain classes of stochastic parabolic equations with Robin boundary conditions has been addressed in [10, 43]. For null controllability results concerning stochastic semidiscrete parabolic equations, we refer the reader to [48].
@@
Regarding controllability results for stochastic parabolic equations with dynamic boundary conditions, we refer to [7] for forward and backward stochastic parabolic equations involving only zero-order terms, to [6] for a class of one-dimensional stochastic heat equations, to [5] for backward stochastic parabolic equations incorporating both reaction and convection terms, and to [11] for the forward stochastic heat equation with a time–space localized control acting in the drift term of the bulk equation, without additional forcing terms. For controllability results concerning certain coupled forward–backward stochastic parabolic equations, we refer the reader to [29, 38].
@@
Following the classical duality approach, the controllability of linear stochastic partial differential equations, as studied in the aforementioned works, can be established by analyzing an observability problem for the corresponding adjoint equation. This approach relies on unique continuation properties to achieve approximate controllability and on observability inequalities to study null controllability and insensitizing control problems. A fundamental tool for proving such observability results—both for parabolic PDEs and SPDEs—is the use of Carleman estimates, which provide weighted energy estimates with sufficiently large parameters to absorb lower-order terms. Introduced by Carleman in 1939 to establish uniqueness results for second-order elliptic equations (see [13]), these estimates have played a crucial role in the study of control and inverse problems in the deterministic setting (see, e.g., [42]). More recently, in the stochastic setting, Carleman estimates have also become a central tool for studying control theory (see, e.g., [34]) as well as for analyzing inverse problems (see, for instance, [32, 33]).
For insensitizing control problems for certain classes of stochastic parabolic equations, we refer the reader to the following works: the forward stochastic heat equation is studied in [44], backward stochastic heat equations in [29], and forward stochastic degenerate parabolic equations in [30]. Insensitizing control problems for a class of stochastic forward parabolic equations with dynamic boundary conditions are considered in [3]. The stochastic Kuramoto–Sivashinsky equation is treated in [35]. Against this background, the present paper investigates, for the first time, insensitizing control problems for forward anisotropic stochastic reaction–convection–diffusion equations with general dynamic boundary conditions.
@@
To the best of our knowledge, this work is the first to address the null controllability and an insensitizing control problem for forward anisotropic stochastic parabolic equations with general dynamic boundary conditions, including both reaction and convection terms. Furthermore, we provide the explicit estimate (1.15) for the null controllability cost of (CE), expressed in terms of the final control time $T$ and the problem data. The cost of approximate controllability of (CE) can also be considered; however, it is left open in the present paper. For related results on the deterministic heat equation, we refer the reader to [9, 17].
In equation (1.9), the control functions $u$, $v_{1}$, and $v_{2}$ are introduced to influence the evolution of the system and, in particular, its behavior at the final time $T$. The conormal derivative $\partial_{\nu}^{\mathcal{A}}y$ acts as the coupling term between the bulk equation in $Q$ and the surface equation on $\Sigma$. Consequently, the distributed control $u$ directly affects the drift term of the bulk equation on the interior subdomain $G_{0}$, while the controls $v_{1}$ and $v_{2}$ act on the diffusion terms of the bulk and surface equations, respectively, over the entire spatial domain. Through the coupling term $\partial_{\nu}^{\mathcal{A}}y$, the controls $u$, $v_{1}$, and $v_{2}$ also indirectly influence the drift term of the surface equation. This interaction between bulk and boundary dynamics is a key mechanism that allows us to study the null controllability of (CE) and the insensitizing control problem (1.11). By contrast, investigate these control problems using only the control $u$, which acts exclusively on the drift term of the bulk equation, remains a challenging open problem; for partial results in this direction under the additional assumption of space-independent coefficients, we refer the reader to [11, 31].
In equation (1.9), the control functions $u$, $v_{1}$, and $v_{2}$ are introduced to influence the evolution of the system, and in particular its behavior at the final time $T$. The conormal derivative $\partial_{\nu}^{\mathcal{A}}y$ acts as the coupling term between the bulk equation in $Q$ and the surface equation on $\Sigma$. Consequently, the distributed control $u$ directly affects the drift term of the bulk equation on the interior subdomain $G_{0}$, while the controls $v_{1}$ and $v_{2}$ act on the diffusion terms of the bulk and surface equations, respectively, over the entire spatial domain. Through the coupling term $\partial_{\nu}^{\mathcal{A}}y$, the controls $u$, $v_{1}$, and $v_{2}$ also indirectly influence the drift term of the surface equation. This interaction between bulk and boundary dynamics is a key mechanism that allows us to study the null controllability of (CE) and the insensitizing control problem (1.11). By contrast, investigating these control problems using only the control $u$, which acts exclusively on the drift term of the bulk equation, remains a challenging open problem.
The Carleman estimate established in [7, Theorem 4.1] for backward stochastic parabolic equations, obtained via a weighted identity method for (1.2) with drift terms in $\mathbb{L}^{2}$ (that is, $F_{j}={F_{\Gamma}}_{j}=0$ for all $j=1,2,\dots,N$), is not sufficient to address the control problems considered in this paper. In particular, it cannot be applied to the null controllability problem studied in Section 5, nor to the insensitizing control problem analyzed in Section 6. To overcome this limitation, an observability inequality for the associated adjoint problem with drift terms in negative-order Sobolev spaces is required; such an estimate is not provided by [7, Theorem 4.1]. For this reason, we derive and employ the Carleman estimate (1.8) in the present work, which enables us to solve the null controllability problem in Section 5. Furthermore, in Section 6, we reduce the insensitizing control problem (1.11) to a null controllability problem for an optimality cascade forward-backward system. This system involves zeroth-, first-, and second-order coupling terms in the bulk–surface equations. The first- and second-order (weak) coupling terms are of the form
$\nabla\cdot(zB_{1}),\qquad\nabla_{\Gamma}\cdot\bigl(z_{\Gamma}B_{2}-\mathbbm{1}_{\mathcal{O}_{\Gamma}^{2}}\nabla_{\Gamma}y_{\Gamma}\bigr),$
and their treatment requires the use of the Carleman estimate (1.8) to derive a new Carleman estimate for the considered adjoint coupled system (see Proposition 6.2).
@@
In the derivation of the Carleman estimate (1.8) for (1.2), the boundary vector fields
$\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma}=\Bigl(\sum_{k=1}^{N}a^{jk}_{\Gamma}D_{k}z_{\Gamma}\Bigr)_{1\leq j\leq N}\quad\textnormal{and}\quad F_{\Gamma}=\bigl({F_{\Gamma}}_{j}\bigr)_{1\leq j\leq N}$
are assumed to be tangential to the manifold $\Gamma$. When this tangentiality condition is not satisfied, one must instead employ the following generalized tangential divergence formula (see [20, Proposition 5.4.9]):
$\int_{\Gamma}(\nabla_{\Gamma}\cdot f_{\Gamma})\,z_{\Gamma}\,d\sigma=-\int_{\Gamma}f_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}\,d\sigma+\int_{\Gamma}(\nabla_{\Gamma}\cdot\nu)\,(f_{\Gamma}\cdot\nu)\,z_{\Gamma}\,d\sigma.$ (1.19)
The additional term in the right-hand side of (1.19), involving the normal component $f_{\Gamma}\cdot\nu$, generates extra boundary contributions in the Carleman estimate. These terms cannot be controlled by the techniques developed in the present paper and, in particular, we do not know how to absorb them into the left-hand side of the Carleman inequality. As a consequence, the analysis of non-tangential boundary vector fields within the framework of Carleman estimates for equation (1.2) remains unsolved in this paper.
Remark 1.2.
In equation (CE), the controls $v_{1}$ and $v_{2}$ act on the diffusion terms in the whole domain $G$ and on its boundary $\Gamma$, respectively. This allows the controllability results for (CE) to be extended to systems where the diffusion terms depend on the state and its gradients. Specifically, the terms $v_{1}\,dW(t)$ and $v_{2}\,dW(t)$ can be replaced by
$(b_{1}y+\widetilde{B}_{1}\cdot\nabla y+v_{1})\,dW(t)\quad\textnormal{and}\quad(b_{2}y_{\Gamma}+\widetilde{B}_{2}\cdot\nabla_{\Gamma}y_{\Gamma}+v_{2})\,dW(t),$
where $b_{i}$ and $\widetilde{B}_{i}$, $i=1,2$, are bounded, adapted stochastic processes. Indeed, if system (CE) is null controllable, then there exists a control triple $(u,v_{1},v_{2})$ such that
@@
Defining the new control triple
$u^{\star}=u,\qquad v_{1}^{\star}=v_{1}-(b_{1}y+\widetilde{B}_{1}\cdot\nabla y),\qquad v_{2}^{\star}=v_{2}-(b_{2}y_{\Gamma}+\widetilde{B}_{2}\cdot\nabla_{\Gamma}y_{\Gamma}),$
one obtains null controllability for the extended system
$\begin{cases}dy-\nabla\cdot(\mathcal{A}\nabla y)\,dt=(a_{1}y+B_{1}\cdot\nabla y+\mathbbm{1}_{G_{0}}u^{\star})\,dt+(b_{1}y+\widetilde{B}_{1}\cdot\nabla y+v_{1}^{\star})\,dW(t),&\textnormal{in }Q,\\[2.84526pt] dy_{\Gamma}-\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma})\,dt+\partial_{\nu}^{\mathcal{A}}y\,dt=(a_{2}y_{\Gamma}+B_{2}\cdot\nabla_{\Gamma}y_{\Gamma})\,dt+(b_{2}y_{\Gamma}+\widetilde{B}_{2}\cdot\nabla_{\Gamma}y_{\Gamma}+v_{2}^{\star})\,dW(t),&\textnormal{on }\Sigma,\\ y_{\Gamma}=y|_{\Gamma},&\textnormal{on }\Sigma,\\ (y,y_{\Gamma})|_{t=0}=(y_{0},y_{\Gamma,0}),&\textnormal{in }G\times\Gamma.\end{cases}$